FGC Software Design Review - LHC commissioning_4_ by hcj

VIEWS: 11 PAGES: 11

									        The programme – as presented yesterday




15/02/2013             LHC 8:30 meeting
        Programme last evening




15/02/2013             LHC 8:30 meeting
        Programme this morning




15/02/2013             LHC 8:30 meeting
        Thursday 14th February
     n 07:24 Last dump from Stable Beams for a long time…
             ¨ Operator dump!
     n 08:30 At injection, start of DS Quench Tests.
     n 13:19 Dump ramp 1 of collimation quench test at dispersion suppressor,
       slow losses,
             ¨ 3 bunches for ADT tests, with asynch dump
             ¨ Asynch dump OK
     n 15:09 Dump ramp 2 of collimation quench test due to interlock from
       collimators when moving to relaxed settings
             ¨ Several source, open TCDQ energy threshold
     n Access for installing scope Q6 quench test at injection
             ¨ Took down the sector, not foreseen by some
             ¨ Pre-cycle etc.
     n 20:21 Dump ramp 3 of DS quench test, 144 bunches, about 500 kW
       losses, dumped by BLMs without quench
             ¨ Increase BLM limits


15/02/2013                                 LHC 8:30 meeting
        First ‘real’ DS test ramp with 144 bunches
    n The beam was dumped 3 seconds before the excitation
      strength started to increase to provide constant loss rate.


    Blue - excitation strength
    Green - intensity
    Red - excitation running




15/02/2013                       LHC 8:30 meeting
        By the collimation team:
    n In more detail
             ¨ dumped on the RS10 for the BLMQI.09L7.B2I10_MQ; for this
               magnet the MF was still at 0.1.
             ¨ We achieved a peak loss rate of 8-8.5e11p/s over about 2 s
               (exceeding what was achieved in 2011).
             ¨ Other faster RS's achieved up to 85% of dump levels (TCPs,
               TCLAs).
             ¨ We achieved (without quenching) losses up to 1.5 times larger
               than the assumed quench limit (first time for proton losses).
    n We decided to proceed as it follows:
             ¨ Increase MF on a selection of DS magnets to 0.5.
             ¨ Change the master thresholds by a factor 2.5 for the collimators
               that are close to dump limit for fast RS.
             ¨ In the meantime, deploy the firmware change for the ADT to
               allow excitation on 2 trains.


15/02/2013                             LHC 8:30 meeting
        Continue DS quench test

    n 00:07 Dump again on BLMs (no quench)




15/02/2013                LHC 8:30 meeting
        Summary of second attempt to quench.
    n Achieved peak loss rates of 640 kW, corresponding to
      up to 1e12p/s.
    n Dumped on RS10 on the BLMQI.08L7.B2I10_MQ. We
      reached 3 times the quench limit in RS10, without
      quenching.
             ¨ This monitor has MF=1 so we can go above that levels only by
               changing the MT.
    n Next attempt:
             ¨ Increase the intensity by a factor ~1.5 (144+72b)
             ¨ changing the limiting Q8 monitor by a factor 2.5 by changing the
               MT (in RS09 and above).
             ¨ Other changes MF's in the DS, listed




15/02/2013                             LHC 8:30 meeting
        DS quench test with B2 (144+72 bunches)

    n Dump at 3:15
    n We achieved peak losses up to 1MW without
      quenching. Peak loss rate: about 1.65e12p/s
    n This time we dumped on a TCLA collimators (thresholds
      were at the electronics maximum)
             ¨ RS08
    n The Q8 reached for RS09 a factor 3.4 above the
      assumed quench limit.
    n The primary collimators experienced a temperature
      increase up to 10 deg.




15/02/2013                  LHC 8:30 meeting
        Quench test at injection Q6.L8

    n 05:46 It did quench
             ¨ As we know we can with injecting beam
    n TCLIB closed
    n Some iteration on current and injected intensity
    n Quenched on
             ¨ RQ6.L8 at 2500 A
             ¨ One shot at 6.54E10
    n However, not sure if it really quenches or just triggered
      the QPS
             ¨ “RQ6.L8 trip analysis...it looks a bit quick to be a quench”
    n Presently
             ¨ Access for scope installation ADT fast losses quench test
             ¨ Access for IPQ threshold adjustment

15/02/2013                              LHC 8:30 meeting
        The Plan




15/02/2013         LHC 8:30 meeting

								
To top